# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v3

@@ Appendix C Notation and Basic Lemmas
By the law of large numbers, $\lim_{m\to\infty}\hat{\mathcal{T}}_{k,m}=\hat{\mathcal{T}}_{k}$, where the error decays in $O(1/\sqrt{m})$ by the Chernoff bound. Also, $\hat{\mathcal{T}}_{n}\coloneq\mathcal{T}$. Further, Lemma C.4 is independent of the choice of $k$. Therefore, for $k=n$, this implies an identical bound on $Q^{t}$. An identical argument implies the same bound on $\hat{Q}_{k,m}^{t}$.
$\mathcal{T}$ satisfies a $\gamma$-contractive property under the infinity norm [Watkins and Dayan, 1992]. We similarly show that $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k,m}$ satisfy a $\gamma$-contractive property under infinity norm in Lemmas C.6 and C.7.
@@
Similarly to Lemma C.6, suppose we apply $\hat{\mathcal{T}}_{k,m}$ to $\hat{Q}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$ and $\hat{Q}_{k,m}^{\prime}(s_{g},F_{z_{\Delta}},a_{g})$. Then:
@@
The $\gamma$-contractivity of $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k,m}$ attracts the trajectory between two $\hat{Q}_{k}$ and $\hat{Q}_{k,m}$ functions on the same state-action tuple by $\gamma$ at each step. Repeatedly applying the Bellman operators produces a unique fixed-point from the Banach fixed-point theorem which we introduce in Definitions C.9 and C.10.
@@
Further, noting that $\hat{Q}_{k}^{0}=\hat{Q}_{k,m}^{0}:=0$, $\|\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{\tilde{r}}{1-\gamma}$, and $\|\hat{Q}_{k,m}^{\mathrm{est}}\|_{\infty}\leq\frac{\tilde{r}}{1-\gamma}$ from Lemma C.4:
@@
Corollary C.11 characterizes the error decay between $\hat{Q}_{k}^{T}$ and $\hat{Q}_{k}^{*}$ and shows that it decays exponentially in the number of Bellman iterations by a $\gamma^{T}$ multiplicative factor.
@@
Algorithm 3 gives a stable implementation of Algorithm 1 with learning rates $\{\eta_{t}\}_{t\in[T]}$. Algorithm 3 is provably numerical stable under fixed-point arithmetic [Anand et al., 2024, 2025]. $\hat{Q}_{k,m}^{t}$ in Algorithm 3 is $\gamma$-contractive as in Lemma C.6, given an appropriately conditioned sequence of learning rates $\eta_{t}$:
Algorithm 3 Stable (Practical) Implementation of Algorithm 1: SUBSAMPLE-Q: Learning
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Hyperparameter $k\in[n]$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$. Learning rate sequence $\{\eta_{t}\}_{t\in[T]}$ where $\eta_{t}\in(0,1]$.
@@
Furthermore, finite-time guarantees with the learning rate and sample complexity have been shown in Chen and Maguluri [2022b], which when adapted to our $\hat{Q}_{k,m}$ framework in Algorithm 3 yields:
@@ Appendix D Proof Sketch
This section details an outline for the proof of Theorem 4.4, as well as some key ideas. At a high level, our SUBSAMPLE-MFQ framework recovers exact mean-field $Q$ learning and traditional value iteration when $k=n$ and as $m\to\infty$. Further, as $k\!\to\!n$, $\hat{Q}_{k}^{*}$ should intuitively get closer to $Q^{*}$ from which the optimal policy is derived. Thus, the proof is divided into three major steps: firstly, we prove a Lipschitz continuity bound between $\hat{Q}_{k}^{*}$ and $\hat{Q}_{n}^{*}$ in terms of the total variation (TV) distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Next, we bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Finally, we bound the value differences between ${\pi}_{k}^{\mathrm{est}}$ and $\pi^{*}$ by bounding $Q^{*}(s,\pi^{*}(s))-Q^{*}(s,{\pi}_{k}^{\mathrm{est}}(s))$ and then using the performance difference lemma from Kakade and Langford [2002].
@@
We defer the proof of Theorem D.1 to Appendix E. See Figure 4 for a comparison between the $\hat{Q}_{k}^{*}$ learning and estimation process, and the exact ${Q}$-learning framework.
@@
Figure 4: Flow of the algorithm and relevant analyses in learning $Q^{*}$. Here, (1) follows by performing Algorithm 1 (SUBSAMPLE-MFQ: Learning) on $\hat{Q}_{k,m}^{0}$. (2) follows from Lemma 4.3. (3) follows from the Lipschitz continuity and total variation distance bounds in Theorems D.1 and D.2. Finally, (4) follows from noting that $\hat{Q}_{n}^{*}=Q^{*}$.
@@
We bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$, where $\Delta\!\in\!\mathcal{U}\binom{[n]}{k}$. This task is equivalent to bounding the discrepancy between the empirical distribution and the distribution of the underlying finite population. When each $i\!\in\!\Delta$ is uniformly sampled without replacement, we use Lemma F.3 from Anand and Qu [2024] which generalizes the Dvoretzky-Kiefer-Wolfowitz (DKW) concentration inequality for empirical distribution functions.
@@
Then, by Theorem D.2 and the definition of total variation distance from Section 2, we have that for $\delta\in(0,1]$, with probability at least $1-\delta$,
@@
We then apply this result to our MARL setting by studying the rate of decay of the objective function between the learned policy ${\pi}_{k}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$ (Theorem 4.4).
@@
Therefore, we formulate and prove Theorem D.3 which yields a probabilistic bound on this difference, where the randomness is over the choice of $\Delta\in\binom{[n]}{k}$:
@@
We defer the proof of Theorem D.3 and finding optimal value of $\delta$ to Lemma G.8 in the Appendix. Using Theorem D.3 and the performance difference lemma leads to Theorem 4.4.
@@ Appendix E Proof of Lipschitz-continuity Bound
This section proves a Lipschitz-continuity bound between $\hat{Q}_{k}^{*}$ and $Q^{*}$ and includes a framework to compare $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}^{*}_{k}(s_{g},s_{\Delta},a_{g})$ and $Q^{*}(s,a_{g})$ in Lemma E.12.
@@
The following definitions will be relevant to the proof of Theorem E.3.
@@
In the first inequality, we rewrite the expectations over the states as the expectation over the joint transition probabilities. The second inequality then follows from Lemma E.14. To apply it to Lemma E.14, we conflate the joint expectation over $(s_{g},s_{\Delta\cup\Delta^{\prime}})$ and reduce it back to the original form of its expectation. Finally, the third inequality follows from Lemma E.5.
@@
For all $T\in\mathbb{N}$, for any $a_{g}\in\mathcal{A}_{g},s_{g}\in\mathcal{S}_{g},s_{\Delta}\in\mathcal{S}_{l}^{k},a_{\Delta}\in\mathcal{A}_{l}^{k},a^{\prime}_{\Delta}\in\mathcal{A}_{l}^{k}$, and for all joint stochastic kernels $\mathcal{J}_{k}$ as defined in Definition E.4:
@@
The first equality follows from noting that $f(x)=\sum_{x^{\prime}\in\mathcal{X}}f(x^{\prime})\mathbbm{1}\{x=x^{\prime}\}$ and from Fubini-Tonelli’s inequality which allows us to swap the order of summations as the summand is finite. The second equality uses the law of total expectation. The first inequality uses Jensen’s inequality and the triangle inequality. The second inequality uses $\|\tilde{r}_{l}(\cdot,\cdot)\|_{\infty}\leq\|r_{l}(\cdot,\cdot)\|_{\infty}$ which holds as $\|{r}_{l}\|_{\infty}$ is the infinite-norm of the local reward functions and is therefore atleast as large any other element in the image of $r_{l}$. The third inequality follows from the definition of total variation distance, and the final inequality follows from Lemma E.7. This proves the base case.
@@
[…]
@@
[…]
@@
The first equality rewrites the equations with their respective maximizing actions. The first inequality upper-bounds this difference by allowing all terms to share the common action $\tilde{a}$. Using the Bellman equation, the second equality expands $\hat{Q}_{k}^{T+1}$ and $\hat{Q}_{k^{\prime}}^{T+1}$. The second inequality follows from the triangle inequality. The third equality follows by subtracting the common $r_{g}(s_{g}^{\prime},\tilde{a_{g}})$ terms from the reward and noting that the two expectation terms $\mathbb{E}_{s_{g}^{\prime},s_{\Delta}^{\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},\tilde{a}_{g},s_{\Delta},\tilde{a}_{\Delta})}$ can be combined into a single expectation $\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime}\sim\tilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}$.
In the special case where $a_{\Delta}=\tilde{a}_{\Delta}$, we derive a closed form expression for $\tilde{\mathcal{J}}_{k}$ in Lemma E.11. To justify the third inequality, the second term follows from the induction hypothesis and the first term follows from the following derivation:
@@
The above derivation follows from the same argument as in the base case where $r_{l}^{\tilde{a}}(s_{i}^{\prime},s_{g}^{\prime}):=r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})$ for any $i\in\Delta\cup\Delta^{\prime}$. Similarly, if (I) $<$ (II), an analogous argument that replaces $\tilde{a}_{\Delta}$ with $\hat{a}_{\Delta}$ yields the same result.
@@
Given a joint transition probability function $\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}$ as defined in Definition E.4, we can recover the transition function for a single agent $i\in\Delta\cup\Delta^{\prime}$ given by $\mathcal{J}_{1}$ using the law of total probability and the conditional independence between $s_{i}$ and $s_{g}\cup s_{[n]\setminus i}$ in Equation 26. This characterization is crucial in Lemma E.7 and Lemma E.5:
@@
Here, we use a conditional independence property proven in Lemma E.10.
@@
The second equality uses Lemma E.9. The third equality uses the property that each local agent can only have one state/action pair, and the fourth equality vectorizes the indicator variables. Then,
@@
Remark: Note that TV-distance is not an $f$-divergence, and therefore it does not generally satisfy the data processing inequality. Under this lens, Lemma E.7 is novel since it saturates an analogous data-processing relation for the TV distance in our multi-agent setting.
@@
The second equality follows from the conditional independence of $s_{i}^{\prime}$ from $s_{\Delta\cup\Delta^{\prime}\setminus i}$ from Lemma E.10, and the final equality uses the fact that the expectation of an indicator random variable is the probability distribution function of the random variable. ∎
@@
The $Q^{*}$ function is the average value of each of the (${n\choose k}$ sub-sampled $\hat{Q}^{*}_{k}$-functions.
@@
[…]
@@
$\bigg{\|}v-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}v^{\prime}\bigg{\|}_{\infty}$ $\geq\bigg{|}\|v\|_{\infty}-\bigg{\|}\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}v^{\prime}\bigg{\|}_{\infty}\bigg{|}$
@@
The first inequality follows from the $\gamma$-contraction property of the update procedure, and the ensuing equality follows from our bound on the maximum possible value of $Q$ from Lemma C.4 and noting that $\hat{Q}_{k}^{0}:=0$.
@@
$=\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\hat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\tilde{\theta}^{*})$

### v4

@@ Appendix C Notation and Basic Lemmas
By the law of large numbers, $\lim_{m\to\infty}\hat{\mathcal{T}}_{k,m}=\hat{\mathcal{T}}_{k}$, where the error decays in $O(1/\sqrt{m})$ by the Chernoff bound. Also, $\hat{\mathcal{T}}_{n}\coloneq\mathcal{T}$. Further, Lemma˜C.4 is independent of the choice of $k$. Therefore, for $k=n$, this implies an identical bound on $Q^{t}$. An identical argument implies the same bound on $\hat{Q}_{k,m}^{t}$.
$\mathcal{T}$ satisfies a $\gamma$-contractive property under the infinity norm [Watkins and Dayan, 1992]. We similarly show that $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k,m}$ satisfy a $\gamma$-contractive property under infinity norm in Lemmas˜C.6 and C.7.
@@
Similarly to Lemma˜C.6, suppose we apply $\hat{\mathcal{T}}_{k,m}$ to $\hat{Q}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$ and $\hat{Q}_{k,m}^{\prime}(s_{g},F_{z_{\Delta}},a_{g})$. Then:
@@
The $\gamma$-contractivity of $\hat{\mathcal{T}}_{k}$ and $\hat{\mathcal{T}}_{k,m}$ attracts the trajectory between two $\hat{Q}_{k}$ and $\hat{Q}_{k,m}$ functions on the same state-action tuple by $\gamma$ at each step. Repeatedly applying the Bellman operators produces a unique fixed-point from the Banach fixed-point theorem which we introduce in Definitions˜C.9 and C.10.
@@
Further, noting that $\hat{Q}_{k}^{0}=\hat{Q}_{k,m}^{0}:=0$, $\|\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{\tilde{r}}{1-\gamma}$, and $\|\hat{Q}_{k,m}^{\mathrm{est}}\|_{\infty}\leq\frac{\tilde{r}}{1-\gamma}$ from Lemma˜C.4:
@@
Corollary˜C.11 characterizes the error decay between $\hat{Q}_{k}^{T}$ and $\hat{Q}_{k}^{*}$ and shows that it decays exponentially in the number of Bellman iterations by a $\gamma^{T}$ multiplicative factor.
@@
Algorithm˜3 gives a stable implementation of Algorithm˜1 with learning rates $\{\eta_{t}\}_{t\in[T]}$. Algorithm˜3 is provably numerical stable under fixed-point arithmetic [Anand et al., 2024, 2025]. $\hat{Q}_{k,m}^{t}$ in Algorithm˜3 is $\gamma$-contractive as in Lemma˜C.6, given an appropriately conditioned sequence of learning rates $\eta_{t}$:
Algorithm 3 Stable (Practical) Implementation of Algorithm˜1: SUBSAMPLE-Q: Learning
0: A multi-agent system as described in Section˜2. Parameter $T$ for the number of iterations in the initial value iteration step. Hyperparameter $k\in[n]$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$. Learning rate sequence $\{\eta_{t}\}_{t\in[T]}$ where $\eta_{t}\in(0,1]$.
@@
Furthermore, finite-time guarantees with the learning rate and sample complexity have been shown in Chen and Maguluri [2022b], which when adapted to our $\hat{Q}_{k,m}$ framework in Algorithm˜3 yields:
@@ Appendix D Proof Sketch
This section details an outline for the proof of Theorem˜4.4, as well as some key ideas. At a high level, our SUBSAMPLE-MFQ framework recovers exact mean-field $Q$ learning and traditional value iteration when $k=n$ and as $m\to\infty$. Further, as $k\!\to\!n$, $\hat{Q}_{k}^{*}$ should intuitively get closer to $Q^{*}$ from which the optimal policy is derived. Thus, the proof is divided into three major steps: firstly, we prove a Lipschitz continuity bound between $\hat{Q}_{k}^{*}$ and $\hat{Q}_{n}^{*}$ in terms of the total variation (TV) distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Next, we bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Finally, we bound the value differences between ${\pi}_{k}^{\mathrm{est}}$ and $\pi^{*}$ by bounding $Q^{*}(s,\pi^{*}(s))-Q^{*}(s,{\pi}_{k}^{\mathrm{est}}(s))$ and then using the performance difference lemma from Kakade and Langford [2002].
@@
We defer the proof of Theorem˜D.1 to Appendix E. See Figure˜4 for a comparison between the $\hat{Q}_{k}^{*}$ learning and estimation process, and the exact Q{Q}-learning framework.
@@
Figure 4: Flow of the algorithm and relevant analyses in learning $Q^{*}$. Here, (1) follows by performing Algorithm˜1 (SUBSAMPLE-MFQ: Learning) on $\hat{Q}_{k,m}^{0}$. (2) follows from Lemma˜4.3. (3) follows from the Lipschitz continuity and total variation distance bounds in Theorems˜D.1 and D.2. Finally, (4) follows from noting that $\hat{Q}_{n}^{*}=Q^{*}$.
@@
We bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$, where $\Delta\!\in\!\mathcal{U}\binom{[n]}{k}$. This task is equivalent to bounding the discrepancy between the empirical distribution and the distribution of the underlying finite population. When each $i\!\in\!\Delta$ is uniformly sampled without replacement, we use Lemma˜F.3 from Anand and Qu [2024] which generalizes the Dvoretzky-Kiefer-Wolfowitz (DKW) concentration inequality for empirical distribution functions.
@@
Then, by Theorem˜D.2 and the definition of total variation distance from Section˜2, we have that for $\delta\in(0,1]$, with probability at least $1-\delta$,
@@
We then apply this result to our MARL setting by studying the rate of decay of the objective function between the learned policy ${\pi}_{k}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$ (Theorem˜4.4).
@@
Therefore, we formulate and prove Theorem˜D.3 which yields a probabilistic bound on this difference, where the randomness is over the choice of $\Delta\in\binom{[n]}{k}$:
@@
We defer the proof of Theorem˜D.3 and finding optimal value of $\delta$ to Lemma˜G.8 in the Appendix. Using Theorem˜D.3 and the performance difference lemma leads to Theorem˜4.4.
@@ Appendix E Proof of Lipschitz-continuity Bound
This section proves a Lipschitz-continuity bound between $\hat{Q}_{k}^{*}$ and $Q^{*}$ and includes a framework to compare $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}^{*}_{k}(s_{g},s_{\Delta},a_{g})$ and $Q^{*}(s,a_{g})$ in Lemma˜E.12.
@@
The following definitions will be relevant to the proof of Theorem˜E.3.
@@
In the first inequality, we rewrite the expectations over the states as the expectation over the joint transition probabilities. The second inequality then follows from Lemma˜E.14. To apply it to Lemma˜E.14, we conflate the joint expectation over $(s_{g},s_{\Delta\cup\Delta^{\prime}})$ and reduce it back to the original form of its expectation. Finally, the third inequality follows from Lemma˜E.5.
@@
For all $T\in\mathbb{N}$, for any $a_{g}\in\mathcal{A}_{g},s_{g}\in\mathcal{S}_{g},s_{\Delta}\in\mathcal{S}_{l}^{k},a_{\Delta}\in\mathcal{A}_{l}^{k},a^{\prime}_{\Delta}\in\mathcal{A}_{l}^{k}$, and for all joint stochastic kernels $\mathcal{J}_{k}$ as defined in Definition˜E.4:
@@
The first equality follows from noting that $f(x)=\sum_{x^{\prime}\in\mathcal{X}}f(x^{\prime})\mathbbm{1}\{x=x^{\prime}\}$ and from Fubini-Tonelli’s inequality which allows us to swap the order of summations as the summand is finite. The second equality uses the law of total expectation. The first inequality uses Jensen’s inequality and the triangle inequality. The second inequality uses $\|\tilde{r}_{l}(\cdot,\cdot)\|_{\infty}\leq\|r_{l}(\cdot,\cdot)\|_{\infty}$ which holds as $\|{r}_{l}\|_{\infty}$ is the infinite-norm of the local reward functions and is therefore atleast as large any other element in the image of $r_{l}$. The third inequality follows from the definition of total variation distance, and the final inequality follows from Lemma˜E.7. This proves the base case.
@@
$\quad\quad\quad\quad\quad\quad\quad\quad-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\hat{Q}_{k^{\prime}}^{T+1}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg|=\text{(I) - \text{(II)}},$
@@
$\text{(II)}=$ $\bigg|\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime}}\bigg[r_{\Delta^{\prime}}(s_{g}^{\prime},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})$
@@
[…]
In the special case where $a_{\Delta}=\tilde{a}_{\Delta}$, we derive a closed form expression for $\tilde{\mathcal{J}}_{k}$ in Lemma˜E.11. To justify the third inequality, the second term follows from the induction hypothesis and the first term follows from the following derivation:
@@
The above derivation follows from the same argument as in the base case where $r_{l}^{\tilde{a}}(s_{i}^{\prime},s_{g}^{\prime}):=r_{l}(s_{i}^{\prime},s_{g}^{\prime},\tilde{a}_{i})$ for any $i\in\Delta\cup\Delta^{\prime}$. Similarly, if (I) << (II), an analogous argument that replaces $\tilde{a}_{\Delta}$ with $\hat{a}_{\Delta}$ yields the same result.
@@
Given a joint transition probability function $\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}$ as defined in Definition˜E.4, we can recover the transition function for a single agent $i\in\Delta\cup\Delta^{\prime}$ given by $\mathcal{J}_{1}$ using the law of total probability and the conditional independence between $s_{i}$ and $s_{g}\cup s_{[n]\setminus i}$ in Equation˜28. This characterization is crucial in Lemma˜E.7 and Lemma˜E.5:
@@
Here, we use a conditional independence property proven in Lemma˜E.10.
@@
The second equality uses Lemma˜E.9. The third equality uses the property that each local agent can only have one state/action pair, and the fourth equality vectorizes the indicator variables. Then,
@@
Remark: We show an analog of the data processing inequality. Under this lens, Lemma˜E.7 saturates the data-processing relation for the TV distance in our multi-agent setting.
@@
The second equality follows from the conditional independence of $s_{i}^{\prime}$ from $s_{\Delta\cup\Delta^{\prime}\setminus i}$ from Lemma˜E.10, and the final equality uses the fact that the expectation of an indicator random variable is the probability distribution function of the random variable. ∎
@@
The $Q^{*}$ function is the average value of each of the $n\choose k$} sub-sampled $\hat{Q}^{*}_{k}$-functions.
@@
In the second equality, we reparameterized the sum to count the number of times each $r_{l}(s_{i},s_{g})$ was added for each $i\in\Delta$. To do this, we count the number of (k1)(k-1) other agents that could form a $k$-tuple with agent $i$, and there are (n1))(n-1)) candidates from which we chooses the (k1)(k-1) agents. In the last equality, we expanded and simplified the binomial coefficients.
@@
[…]
@@
The first inequality follows from the $\gamma$-contraction property of the update procedure, and the ensuing equality follows from our bound on the maximum possible value of $Q$ from Lemma˜C.4 and noting that $\hat{Q}_{k}^{0}:=0$.
@@
[…]